Please help me solve this problem.
.25x+.5(4-x)=.10(10)

Answer by checkley77(12844)

.25x+.5(4-x)=.10(10)
.25x+2-.5x=1
-.25x=1-2
-.25x=-1
x=-1/-.25
x=4 ans.
Proof:
.25*4+.5(4-4)=.1*10
1+5*0=1
1=1